# Towards gravimetry enhancement with squeezed states

### v1

@@ Abstract
We investigate the estimation sensitivity of gravitational acceleration using squeezed probe states within a quantum metrology framework. In particular, we analyze how the squeezing phase, beyond its amplitudes, of the probes affects the attainable precision. We find that probes squeezed along the canonical phase-space quadrature can fail to achieve a quantum Fisher information (QFI) surpassing the shot-noise limit, regardless of the interaction time with the gravitational field. In contrast, position-momentum correlated input states with the squeezing amplitude can overcome this limit. Furthermore, we show that optimal sensitivity is attained through projective momentum measurements combined with a time-dependent adjustment of the squeezing phase. Our results are important to highlight the fundamental role of phase-engineered squeezing in experimental gravimetry protocols.
@@
Introduction. The pursuit of novel protocols towards high-precision measurements of gravitational acceleration is crucial in the field of gravimetry [1, 2, 3, 4]. Experimentally, significant advances have been made in gravimeters employing several quantum techniques, including atomic interferometry [5, 6, 7], Bose–Einstein condensate (BEC) [4, 8, 9], ultracold neutrons [10], superconducting systems [11], and optomechanics systems  [12, 13]. Theoretically, improvements in the sensing of gravitational fields are pivotal in fundamental tests of general relativity theory [14, 15, 16, 17, 18] and also in the ongoing challenges to unify quantum mechanics and general relativity [19, 20, 21, 22]. Furthermore, from a technological perspective, ultrasensitive gravimeters are essential for developments in geological surveys and navigation, for instance [23, 24, 25].
@@
Accurate estimation of physical quantities is crucial for theoretical developments and experimental breakthroughs. In estimation theory, the Cramér-Rao inequality, $\text{Var}(\hat{\alpha})\geq[n\mathcal{I}_{\alpha}(x)]^{-1/2}$, establishes an essential lower bound on the sensitivity (i.e., the standard deviation) with which an unknown parameter, $\alpha$, encoded in a conditional probability $p(x|\alpha)$ of observing the random value $x$, can be estimated [26, 27]. Here, $\hat{\alpha}$ denotes an unbiased estimator that maps the outcome $x$ to an estimate of the parameter $\alpha$, $n$ is the number of independent experiment trials, and $\mathcal{I}_{\alpha}(x)$ is the classical Fisher information (CFI). The CFI quantifies how changes in the parameter value $\alpha$ affect the probability distribution of the measured quantity $x$, through a particular choice of measurement strategy or a set of positive operator-valued measurements (POVM) [28, 29]. Given its dependence on the specific measurement strategy, the CFI can be optimized for all possible POVM [26, 28, 29], leading to an ultimate upper limit for the CFI, namely Quantum Fisher information (QFI), $\mathcal{F}_{\alpha}=\text{Tr}[\hat{\rho}_{\alpha}\hat{L}_{\alpha}]$, where $\hat{L}_{\alpha}$ is the symmetric logarithmic derivative defined as $\partial_{\alpha}\hat{\rho}_{\alpha}=(\hat{L}_{\alpha}\hat{\rho}_{\alpha}+\hat{\rho}_{\alpha}\hat{L}_{\alpha})/2$ [30, 28, 29]. The quantum parameter estimation theory establishes the QFI as a fundamental precision limit for estimating the parameter $\alpha$, bounding the classical Fisher information, i.e., $\mathcal{F}_{\alpha}\geq\mathcal{I}_{\alpha}$ and giving rise to the quantum Cramér–Rao inequality, which defines the minimal achievable estimation variance employing a quantum state [31, 28, 29, 30, 32].
@@
Quantum metrology aims to explore the features of quantum probes such as entanglement [33, 28, 29], coherence [34], non-classicality [35, 36], contextuality [37], and squeezing [38, 39, 40, 41], to overcome the precision of sensing unknown parameters compared to their classical counterparts [32, 28, 29], leading to a series of experimental advances in this field [42, 43]. In particular, since their original conception [44, 45] and pioneering proposal for use in laser interferometric gravitational wave detectors [38], squeezed states have found applications in a wide range of gravimetry protocols [46]. Despite the recognized advantage of employing squeezed probe states in gravimetry protocols, most applications have mainly focused on the role of the squeezing amplitude or have considered situations where the squeezing phase is aligned with a canonical phase-space quadrature, such as position or momentum.
@@
In this Letter, we investigate the boost in the estimation of the gravitational acceleration, $g$, by considering input squeezed states endowed with a squeezing phase that differs from the canonical quadratures of position and momentum. Specifically, we analyze how the gravitational acceleration is dynamically encoded in both the amplitude and phase of the squeezed probe and how different phases of squeeze affect the QFI. We demonstrate that by choosing an appropriate phase of the squeezed input state, one can create a non-null correlation between the position and momentum quadratures, thereby significantly increasing the QFI and improving the sensitivity in estimating the gravitational acceleration, which overcomes vacuum probe states, regardless of the free-falling time under the gravitational field. Additionally, we show that a projective momentum measurement can saturate the QFI bound and therefore achieve this optimal sensitivity. Notably, standard quadrature squeezing along position or momentum does not always yield optimal results, highlighting the importance of the fine-tuning of the squeezing phase in quantum metrology protocols.
@@
that formally corresponds to a squeezed vacuum state, where $r$ and $\theta$ represent the squeezing amplitude and phase, respectively. $\mathcal{N}_{0}=1/(\sqrt{2\pi}\sigma)$ denotes the normalization constant, $\sigma=\sigma_{0}[\cosh(2r)-\sinh(2r)\cos(2\theta)]^{1/2}$ is the position uncertainty at initial time, $\sigma_{0}$ the standard deviation of the vacuum state, and $\gamma=\sinh(2r)\sin(2\theta)$ quantifies the position-momentum correlation [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The phase $\theta$ sets the quadrature along which squeezing is applied and determines the orientation of the uncertainty ellipse in phase space (See Supplemental Material (SM) [51]). For instance, when $\theta=0$ the squeezing is along the position quadrature, $\sigma=\sigma_{0}/e^{2r}$ (reduction in position uncertainty), and for $\theta=\pi/2$, the squeezing is along the momentum quadrature, $\sigma=\sigma_{0}e^{2r}$ (increasing in position uncertainty). For intermediate values of $\theta$, the squeezed input state (1) exhibits a non-null correlation between the position and momentum quadratures [57]. Recent studies have highlighted the role of such position-momentum correlations as a valuable resource in metrological tasks [58, 59, 60, 61]. Position-momentum correlations are a key resource for universal continuous-variable quantum computation (CVQC) [62], as established by a recent resource theory [63].
@@
where $\tau_{0}=m\sigma_{0}^{2}/\hbar$ is a constant with time dimension. In the particular case where $r=0$ (i.e., $\gamma=0$), the QFI (2) becomes $\mathcal{F}_{g}^{\text{vac}}(\tau)=\tau^{4}/4\sigma^{2}_{0}+4\tau_{0}^{2}\tau^{2}/\sigma^{2}_{0}$, consistent with the result obtained in Ref. [64] for an input vacuum state. Some conclusions can be drawn from Eq. (2). First, the QFI depends on both amplitude and phase of the squeezing. Consequently, it allows for the possibility of setting up different squeezing phases of the input probe, rather than restricting squeezing to the conventional quadratures by fixing $\theta=0$ or $\pi/2$, to enhance the sensitivity of gravitational acceleration compared to the input vacuum state. Moreover, since the energy of the input probe state is independent of $\theta$, possible advantages in the QFI by changing the squeezing phase come without any additional energy cost. Second, the QFI is insensitive to the intensity of the gravitational acceleration, meaning that the capacity to estimate the $g$ does not change with the strength of this parameter. Third, for short ($\tau\ll\tau_{0}$) and long ($\tau\gg\tau_{0}$) interaction times, the QFI (2) evolves into $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\theta)\approx[\sigma^{2}/\sigma_{0}^{2}]\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ and $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\theta)\approx[\sigma_{0}^{2}/\sigma^{2}+\gamma^{2}]\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$, respectively. This characteristic enables fine-tuning the squeezing intensity $r$ and phase $\theta$ of the input probe to enhance beyond that of the vacuum state, thereby increasing the sensitivity to $g$ over specific interaction times.
@@
Figure 1 presents the QFI (2) for estimating the gravitational acceleration parameter $g$ as a function of the dimensionless interaction time $\tau/\tau_{0}$, considering different input probe states. In all plots, we set the following parameters $m=1$, $\sigma_{0}=1$, $\hbar=1$ to focus only on the general trends and features of the dynamics. Panels (a-c) correspond to the dynamic of $\mathcal{F}_{g}(\tau,r,\theta)$ for short interaction times, whereas panels (d-f) illustrate the behavior for long times. The curves represent the squeezing intensities $r=0.4$ (solid black line), $r=0.5$ (dashed red line), $r=0.6$ (dotted blue line), for distinct squeezing phases of the input state (1), and the vacuum state $r=0$ (dot-dashed gray line). For the input states squeezed in momentum ($\theta=\pi/2$), the QFI surpasses that of the vacuum probe only in the short-time regime, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\pi/2)\approx e^{2r}\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [shaded area in Fig. 1(a)], the input vacuum state being more sensitive in estimating the gravitational acceleration $g$ for long times, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\pi/2)\approx e^{-2r}\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(d)]. Conversely, for probes squeezed in position ($\theta=0$), the QFI enhancement over the input vacuum state occurs exclusively in the long time limit, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,0)\approx e^{2r}\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [shaded area in Fig. 1(e)], with the input vacuum state more sensitive for short times, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,0)\approx e^{-2r}\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(b)]. Interestingly, probe states initially endowed with position-momentum correlations ($\theta=\pi/4$) achieve a QFI greater than that of the vacuum state in both the short, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\pi/4)\approx\cosh(2r)\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(c)], and long, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\pi/4)\approx(\sinh^{2}2r+\operatorname{sech}2r)\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(f)], interaction-time regimes, respectively. This result highlights the fact that the quantum advantage in estimating gravitational acceleration using squeezed probes initialized in the canonical position ($\theta=0$) and momentum ($\theta=\pi/2$) direction is limited to a certain time domain. In contrast, squeezed probes endowed with initial position-momentum correlation can achieve a consistent improvement in the QFI over the vacuum state, regardless of the interaction time.
@@ 0 $e^{-2r}$ $e^{2r}$
Despite our choice of $\theta=\pi/4$ in Fig. 2(a), there are other position-momentum correlated state with a QFI surpassing that of the vacuum state. Figure 2(b) displays a contour plot of $\mathcal{Q}$ as a function of both time $\tau$ and the squeezing phase $\theta$, allowing us to explore the continuous variation of metrological performance in different squeezing orientations. In particular, the results reveal that the enhanced sensitivity for all times ($\mathcal{Q}>1$) is not always limited to the specific case $\theta=\pi/4$, but can be achieved across a wide range of angles, highlights that optimal metrological gain for estimation of gravitational acceleration does not depend solely on the squeezing strength but critically on the phase-space orientation of the squeezed quadratures.
@@
Measurement schemes. Although the QFI establishes the ultimate precision bound for an unbiased estimator, $\hat{\alpha}$, the optimal measurement required to achieve this bound is often difficult or impractical [28, 65, 66]. This is because it requires a projective measurement associated with POVM elements given by projectors onto the eigenstates of the logarithmic derivative operator $\hat{L}_{\alpha}$. To identify measurement schemes capable of approaching the bound imposed by the QFI, we focus on strategies that can be readily described within the Gaussian formalism [67]. We consider local Gaussian projective measurements, which can be efficiently modeled by updating the covariance matrix as $\widetilde{\boldsymbol{\sigma}}=\boldsymbol{\sigma}+\boldsymbol{\sigma}_{m}$ [67, 66], where the matrix $\boldsymbol{\sigma}_{m}=\text{Diag}[s,s^{-1}]/2$ characterizes the measurement and is parametrized by $s$. This framework encompasses projective position measurements ($s\rightarrow 0$), projective momentum measurements ($s\rightarrow\infty$), as well as homodyne and heterodyne detection ($s=1$). The exact CFI, $\mathcal{I}_{g}(\tau,r,\theta,s)$, that quantifies the information extracted by a given measurement strategy [67, 66] can be analytically determined in terms of the parameter $s$ (see supplementary material [51]). By scanning the values of $s$, it is possible to identify which measurement strategy is most effective in achieving the ultimate bound, where the CFI approaches the QFI.
@@
Fig. 3: Behavior of the ratio $\mathcal{R}(\tau,r,\theta,s)=\mathcal{I}_{g}(\tau,r,\theta,s)/\mathcal{F}_{g}(\tau,r,\theta)$ between the CFI and the QFI bound for different measurement schemes, as a function of the squeezing phase $\theta$ and the evolution time $\tau$. Here, the squeezing parameter $r=0.5$ is kept fixed across all panels and the measurement setting $s$ varies from one panel to another. Panel (a) shows the case $s\rightarrow\infty$ (projective momentum measurement), (b) corresponds to $s\rightarrow 0$ (projective position measurement), and (c) illustrates the heterodyne scheme with $s=1$. Note that only the momentum measurement strategy can saturate the QFI bound, yielding the maximum possible information $\mathcal{R}=1$, yellow region in Fig. 3(a)]. Yet, the optimal choice of squeezing phase is time-dependent, and squeezing along canonical quadratures, position ($\theta=0$) or momentum ($\theta=\pi/2$), does not always provide the best sensitivity at all times.
@@
In Fig. 3, we show the ratio $\mathcal{R}(\tau,r,\theta,s)=\mathcal{I}_{g}(\tau,r,\theta,s)/\mathcal{F}_{g}(\tau,r,\theta)$ between the CFI, $\mathcal{I}_{g}(\tau,r,\theta,s)$, and the QFI bound, $\mathcal{F}_{g}(\tau,r,\theta)$, for different measurement strategies. The ratio $\mathcal{R}$ quantifies how closely a given measurement approaches the ultimate precision bound set by the QFI. Here, the squeezing parameter $r=0.5$ is kept fixed in all panels, and the measurement setting $s$ varies in each. We analyze how different squeezing phases $\theta$ of the squeezed input state can be adjusted over time to reach the QFI limit. The limiting situation $s\rightarrow\infty$, which denotes an ideal momentum measurement, is shown in Fig. 3(a). This approach can theoretically reach the highest sensitivity allowed by quantum physics in this domain, since the CFI can saturate the QFI bound, i.e., $\mathcal{I}_{g}=\mathcal{F}_{g}$ ($\mathcal{R}=1$) [lighter yellow region, in Fig. 3(a)]. However, this optimality depends on the appropriate choice of the squeezing phase, which changes with time $\tau$. The complementary scenario $s\rightarrow 0$, corresponding to a projective position measurement, is displayed in Fig. 3(b). Unlike the momentum example, the projective position measurement does not saturate the QFI constraint, $\mathcal{R}\approx 0.4$ [yellow region in Fig. 3(b)]. The heterodyne detection scheme with $s=1$ is shown in Figure 3(c). This approach also fails to reach the QFI limit $\mathcal{R}\approx 0.5$, yellow region in Fig. 3(c)], highlighting the importance of searching for appropriate measurement strategies in quantum metrology. Interestingly, optimal sensitivity is not always achieved by canonical quadrature squeezing, i.e., squeezed probe with phase along position ($\theta=0$) or momentum ($\theta=\pi/2$). In addition, the time dependence of the optimal phase $\theta_{\text{opt}}(\tau)$ suggests that adaptive or dynamically controlled squeezed strategies may be needed to approach the quantum limit over different time scales consistently.

### v2

@@ Abstract
We investigate the sensitivity of gravitational acceleration estimation using squeezed probe states in a quantum metrology framework. In particular, we analyze how the squeezing phase, beyond its amplitude, affects the attainable precision. We show that probes squeezed along the canonical phase-space quadratures can surpass the shot-noise limit only in specific time regimes, whereas position–momentum correlated input states can consistently overcome this limit across all interaction times. Furthermore, we demonstrate that optimal sensitivity can be achieved by combining projective momentum measurements with a time-dependent adjustment of the squeezing phase. Our results highlight the fundamental role of phase-engineered squeezing in quantum gravimetry protocols and provide new insights into the design of optimized sensing strategies.
@@
Introduction. The pursuit of novel protocols towards high-precision measurements of gravitational acceleration is crucial in the field of gravimetry [62, 75, 34, 12]. Experimentally, significant advances have been made in gravimeters employing several quantum techniques, including atomic interferometry [57, 4, 3], Bose–Einstein condensate (BEC) [12, 2, 1], ultracold neutrons [39], superconducting systems [50], and optomechanics systems [61, 35]. Theoretically, improvements in the sensing of gravitational fields are pivotal in fundamental tests of general relativity theory [66, 29, 16, 22, 23], ultimate sensing precision in quantum metrology [52], and also in the ongoing challenges to unify quantum mechanics and general relativity [5, 77, 11, 48]. Furthermore, from a technological perspective, ultrasensitive gravimeters are essential for developments in geological surveys and navigation, for instance [18, 15, 73].
@@
Accurate estimation of physical quantities is crucial for theoretical developments and experimental breakthroughs. In estimation theory, the Cramér-Rao inequality, $\text{Var}(\hat{\alpha})\geq[n\mathcal{I}_{\alpha}(x)]^{-1/2}$, establishes an essential lower bound on the sensitivity (i.e., the standard deviation) with which an unknown parameter, $\alpha$, encoded in a conditional probability $p(x|\alpha)$ of observing the random value $x$, can be estimated [37, 38]. Here, $\hat{\alpha}$ denotes an unbiased estimator that maps the outcome $x$ to an estimate of the parameter $\alpha$, $n$ is the number of independent experiment trials, and $\mathcal{I}_{\alpha}(x)$ is the classical Fisher information (CFI). The CFI quantifies how changes in the parameter value $\alpha$ affect the probability distribution of the measured quantity $x$, through a particular choice of measurement strategy or a set of positive operator-valued measurements (POVM) [32, 33]. Given its dependence on the specific measurement strategy, the CFI can be optimized over all possible POVM [37, 32, 33], leading to an ultimate upper limit for the CFI, namely Quantum Fisher information (QFI), $\mathcal{F}_{\alpha}=\text{Tr}[\hat{\rho}_{\alpha}\hat{L}_{\alpha}]$, where $\hat{\rho}_{\alpha}$ is the density operator encoded with the parameter $\alpha$ and $\hat{L}_{\alpha}$ is the symmetric logarithmic derivative defined as $\partial_{\alpha}\hat{\rho}_{\alpha}=(\hat{L}_{\alpha}\hat{\rho}_{\alpha}+\hat{\rho}_{\alpha}\hat{L}_{\alpha})/2$ [9, 32, 33]. The quantum parameter estimation theory establishes the QFI as a fundamental precision limit for estimating the parameter $\alpha$, bounding the classical Fisher information, i.e., $\mathcal{F}_{\alpha}\geq\mathcal{I}_{\alpha}$ and giving rise to the quantum Cramér–Rao inequality, which defines the minimal achievable estimation variance employing a quantum state [10, 32, 33, 9, 41].
@@
Quantum metrology aims to explore the features of quantum probes such as entanglement [59, 32, 33, 14], coherence [63], non-classicality [8, 58, 64], contextuality [40], and squeezing [13, 76, 44, 78], to overcome the precision of sensing unknown parameters compared to their classical counterparts [41, 32, 33], leading to a series of experimental advances in this field [21, 45]. In particular, since their original conception [68, 67] and pioneering proposal for use in laser interferometric gravitational wave detectors [13], squeezed states have found applications in a wide range of gravimetry protocols [65]. Despite the recognized advantage of employing squeezed probe states in gravimetry protocols, most applications have focused mainly on the squeezing amplitude or on situations where the squeezing phase is aligned with a canonical phase-space quadrature, such as position or momentum.
@@
In this Letter, we investigate the boost in the estimation of the gravitational acceleration, $g$, by considering input squeezed states endowed with a squeezing phase that differs from the canonical quadratures of position and momentum. Specifically, we analyze how the gravitational acceleration is dynamically encoded in both the amplitude and phase of the squeezed probe and how different phases of squeeze affect the QFI. We demonstrate that by choosing an appropriate phase of the squeezed input state, one can create a non-zero correlation between the position and momentum quadratures, thereby significantly increasing the QFI and improving the sensitivity in estimating the gravitational acceleration, thereby overcoming vacuum probe states, regardless of the free-falling time under the gravitational field. Additionally, we show that a projective momentum measurement can saturate the QFI bound and therefore achieve this optimal sensitivity. Notably, standard quadrature squeezing along position or momentum does not always yield optimal results, highlighting the importance of the fine-tuning of the squeezing phase in quantum metrology protocols.
@@
that formally corresponds to a squeezed vacuum state characterized by the initial position uncertainty $\sigma=\sigma_{0}[\cosh(2r)-\sinh(2r)\cos(2\theta)]^{1/2}$ and the position-momentum correlation factor $\gamma=\sinh(2r)\sin(2\theta)$ [27, 26, 25, 24, 70, 20, 46, 42, 19, 47]. Here, $\mathcal{N}_{0}=1/(\sqrt{2\pi}\sigma)$ denotes the normalization constant, $r$ and $\theta$ represent the squeezing amplitude and phase, respectively, and $\sigma_{0}$ is the standard deviation of the vacuum state. The phase $\theta$ sets the quadrature along which squeezing is applied and determines the orientation of the uncertainty ellipse in phase space (see Supplemental Material (SM) [70]). For instance, when $\theta=0$ the squeezing is along the position quadrature, $\sigma=\sigma_{0}/e^{2r}$ (reduction in position uncertainty), and for $\theta=\pi/2$, the squeezing is along the momentum quadrature, $\sigma=\sigma_{0}e^{2r}$ (increasing in position uncertainty). For intermediate values of $\theta$, the squeezed input state (1) exhibits a non-null correlation between the position and momentum quadratures [7]. Recent studies have highlighted the role of such position-momentum correlations as a valuable resource in metrological tasks [55, 54, 56, 60]. Position-momentum correlations are a key resource for universal continuous-variable quantum computation (CVQC) [71], as established by a recent resource theory [72].
@@
where $\tau$ is the evolution time and $\tau_{0}=m\sigma_{0}^{2}/\hbar$ is a constant with time dimension. In the particular case where $r=0$ (i.e., $\gamma=0$), the QFI (2) becomes $\mathcal{F}_{g}^{\text{vac}}(\tau)=\tau^{4}/4\sigma^{2}_{0}+4\tau_{0}^{2}\tau^{2}/\sigma^{2}_{0}$, consistent with the result obtained in Ref. [69] for an input vacuum state. Some conclusions can be drawn from Eq. (2). First, the QFI depends on both amplitude and phase of the squeezing. Consequently, it allows setting different squeezing phases for the input probe, rather than restricting squeezing to the conventional quadratures by fixing $\theta=0$ or $\pi/2$, thereby enhancing the sensitivity to gravitational acceleration relative to the input vacuum state. Moreover, since the energy of the input probe state is independent of $\theta$, possible advantages in the QFI by changing the squeezing phase come without any additional energy cost. Second, the QFI is insensitive to the intensity of the gravitational acceleration, meaning that the capacity to estimate the $g$ does not change with the strength of this parameter. Third, for short ($\tau\ll\tau_{0}$) and long ($\tau\gg\tau_{0}$) interaction times limit, the QFI (2) evolves into $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\theta)\approx[\sigma^{2}/\sigma_{0}^{2}]\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ and $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\theta)\approx[\sigma_{0}^{2}/\sigma^{2}+\gamma^{2}]\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$, respectively. This characteristic enables fine-tuning of the squeezing intensity $r$ and phase $\theta$ of the input probe to enhance beyond the vacuum state, thereby increasing the sensitivity to $g$ at specific interaction times.
@@
Figure 1 presents the QFI (2) for estimating the gravitational acceleration parameter $g$ as a function of the dimensionless interaction time $\tau/\tau_{0}$, considering different input probe states. In all plots, we set the following parameters $m=1$, $\sigma_{0}=1$, $\hbar=1$ to focus only on the general trends and features of the dynamics. Panels (a-c) correspond to the dynamic of $\mathcal{F}_{g}(\tau,r,\theta)$ for short interaction times, whereas panels (d-f) illustrate the behavior for long times. The curves represent the squeezing intensities $r=0.4$ (solid black line), $r=0.5$ (dashed red line), $r=0.6$ (dotted blue line), for distinct squeezing phases of the input state (1), and the vacuum state $r=0$ (dot-dashed gray line). For the input states squeezed in momentum ($\theta=\pi/2$), the QFI surpasses that of the vacuum probe only in the short-time regime, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\pi/2)\approx e^{2r}\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [shaded area in Fig. 1(a)], the input vacuum state being more sensitive in estimating the gravitational acceleration $g$ for long times, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\pi/2)\approx e^{-2r}\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(d)]. Conversely, for probes squeezed in position ($\theta=0$), the QFI enhancement over the input vacuum state occurs exclusively in the long time limit, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,0)\approx e^{2r}\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [shaded area in Fig. 1(e)], with the input vacuum state more sensitive for short times, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,0)\approx e^{-2r}\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(b)]. Interestingly, probe states initially endowed with position-momentum correlations ($\theta=\pi/4$) achieve a QFI greater than that of the vacuum state in both the short, $\mathcal{F}_{g}^{\tau\ll\tau_{0}}(\tau,r,\pi/4)\approx\cosh(2r)\mathcal{F}_{g,\tau\ll\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(c)], and long, $\mathcal{F}_{g}^{\tau\gg\tau_{0}}(\tau,r,\pi/4)\approx(\sinh^{2}2r+\operatorname{sech}2r)\mathcal{F}_{g,\tau\gg\tau_{0}}^{\text{vac}}(\tau)$ [Fig. 1(f)], interaction-time regimes, respectively. This result highlights that the quantum advantage in estimating gravitational acceleration using squeezed probes initialized in the canonical position ($\theta=0$) and momentum ($\theta=\pi/2$) directions is limited to a specific time window. In contrast, squeezed probes endowed with initial position-momentum correlation can achieve a consistent improvement in the QFI over the vacuum state, regardless of the interaction time.
Importantly, the QFI does not necessarily increase monotonically with the squeezing strength $r$. This behavior arises from the nontrivial interplay between the hyperbolic functions of $r$ and the trigonometric dependence on the squeezing phase $\theta$ in Eq. 2, which contribute with different weights depending on the interaction time. As a result, increasing $r$ may either enhance or reduce the sensitivity, depending on the squeezing orientation and the time regime. This explains why, in some cases, smaller values of $r$ may outperform larger ones, as observed in Figs. 1(b) and 1(d). This non-monotonic behavior further underscores the central role of the squeezing phase in optimizing metrological performance.
@@ 0 $e^{-2r}$ $e^{2r}$
Despite our choice of $\theta=\pi/4$ in Fig. 2(a), there are other position-momentum correlated states with a QFI surpassing that of the vacuum state. Figure 2(b) displays a contour plot of $\mathcal{Q}$ as a function of both time $\tau$ and the squeezing phase $\theta$, allowing us to explore the continuous variation of metrological performance in different squeezing orientations. In particular, the results reveal that the enhanced sensitivity for all times ($\mathcal{Q}>1$) is not always limited to the specific case $\theta=\pi/4$, but can be achieved across a wide range of angles, highlights that optimal metrological gain for estimation of gravitational acceleration does not depend solely on the squeezing strength but critically on the phase-space orientation of the squeezed quadratures.
@@
Measurement schemes. Although the QFI establishes the ultimate precision bound for an unbiased estimator, $\hat{\alpha}$, the optimal measurement required to achieve this bound is often difficult or impractical [32, 31, 36]. This is because it requires a projective measurement associated with POVM elements given by projectors onto the eigenstates of the logarithmic derivative operator $\hat{L}_{\alpha}$. To identify measurement schemes capable of approaching the bound imposed by the QFI, we focus on strategies that can be readily described within the Gaussian formalism [49]. We consider local Gaussian projective measurements, which can be efficiently modeled by updating the covariance matrix as $\widetilde{\boldsymbol{\sigma}}=\boldsymbol{\sigma}+\boldsymbol{\sigma}_{m}$ [49, 36], where the matrix $\boldsymbol{\sigma}_{m}=\text{Diag}[s,s^{-1}]/2$ characterizes the measurement and is parametrized by $s$. This framework encompasses projective position measurements ($s\rightarrow 0$), projective momentum measurements ($s\rightarrow\infty$), as well as homodyne and heterodyne detection ($s=1$). The exact CFI, $\mathcal{I}_{g}(\tau,r,\theta,s)$, that quantifies the information extracted by a given measurement strategy [49, 36] can be analytically determined in terms of the parameter $s$ (see SM [70]). By scanning the values of $s$, it is possible to identify which measurement strategy is most effective in achieving the ultimate bound, where the CFI approaches the QFI.
@@
Fig. 3: Behavior of the ratio $\mathcal{R}(\tau,r,\theta,s)=\mathcal{I}_{g}(\tau,r,\theta,s)/\mathcal{F}_{g}(\tau,r,\theta)$ between the CFI and the QFI bound for different measurement schemes, as a function of the squeezing phase $\theta$ and the evolution time $\tau$. Here, the squeezing parameter $r=0.5$ is kept fixed across all panels, while the measurement setting $s$ varies from panel to panel. Panel (a) shows the case $s\rightarrow\infty$ (projective momentum measurement), (b) corresponds to $s\rightarrow 0$ (projective position measurement), and (c) illustrates the heterodyne scheme with $s=1$. Note that only the momentum measurement strategy can saturate the QFI bound, yielding the maximum possible information $\mathcal{R}=1$, yellow region in Fig. 3(a)]. Yet, the optimal choice of squeezing phase is time-dependent, and squeezing along canonical quadratures, position ($\theta=0$) or momentum ($\theta=\pi/2$), does not always provide the best sensitivity at all times.
@@
In Fig. 3, we show the ratio $\mathcal{R}(\tau,r,\theta,s)=\mathcal{I}_{g}(\tau,r,\theta,s)/\mathcal{F}_{g}(\tau,r,\theta)$ between the CFI, $\mathcal{I}_{g}(\tau,r,\theta,s)$, and the QFI bound, $\mathcal{F}_{g}(\tau,r,\theta)$, for different measurement strategies. The ratio $\mathcal{R}$ quantifies how closely a given measurement approaches the ultimate precision bound set by the QFI. Here, the squeezing parameter $r=0.5$ is kept fixed across all panels, while the measurement setting $s$ varies between them. We analyze how different squeezing phases $\theta$ of the squeezed input state can be adjusted over time to reach the QFI limit. The limiting situation $s\rightarrow\infty$, which denotes an ideal momentum measurement, is shown in Fig. 3(a). This approach can theoretically reach the highest sensitivity allowed by quantum physics in this domain, since the CFI can saturate the QFI bound, i.e., $\mathcal{I}_{g}=\mathcal{F}_{g}$ ($\mathcal{R}=1$) [lighter yellow region, in Fig. 3(a)]. However, this optimality depends on the appropriate choice of the squeezing phase, which changes with time $\tau$. The complementary scenario $s\rightarrow 0$, corresponding to a projective position measurement, is displayed in Fig. 3(b). Unlike the momentum example, the projective position measurement does not saturate the QFI constraint, $\mathcal{R}\approx 0.4$ [yellow region in Fig. 3(b)]. The heterodyne detection scheme with $s=1$ is shown in Figure 3(c). This approach also fails to reach the QFI limit $\mathcal{R}\approx 0.5$, yellow region in Fig. 3(c)], highlighting the importance of searching for appropriate measurement strategies in quantum metrology. Interestingly, optimal sensitivity is not always achieved by canonical quadrature squeezing, i.e., squeezed probe with phase along position ($\theta=0$) or momentum ($\theta=\pi/2$). In addition, the time dependence of the optimal phase $\theta_{\text{opt}}(\tau)$ suggests that adaptive or dynamically controlled squeezed strategies may be needed to approach the quantum limit over different time scales consistently.
From an experimental perspective, the protocol discussed here is compatible with current quantum gravimetry platforms, such as atom interferometers and atomic fountains [57, 4, 3, 1], as well as other freely falling ultracold atomic systems [39] (see SM [70] for a simulation considering an ensemble of cesium atoms). In these setups, motional Gaussian states can be engineered and manipulated with high precision, including the generation of squeezed states and controlled phase-space rotations [13, 76, 44, 65]. Moreover, projective momentum measurements can be implemented using standard time-of-flight or Doppler-sensitive detection schemes, as routinely employed in ultracold atom experiments [57, 6], while Gaussian measurement strategies, such as homodyne and heterodyne detection, are well established in continuous-variable quantum optics [43, 74]. These considerations indicate that the phase-engineered squeezing strategies proposed here are, in principle, experimentally accessible.